# GraphGSOcc: Semantic and Geometric Graph Transformer for 3D Gaussian Splating-based Occupancy Prediction

### v1

@@ Abstract
Addressing the task of 3D semantic occupancy prediction for autonomous driving, we tackle two key issues in existing 3D Gaussian Splating (3DGS) methods: (1) unified feature aggregation neglecting semantic correlations among similar categories and across regions, and (2) boundary ambiguities caused by the lack of geometric constraints in MLP iterative optimization. We propose the GraphGSOcc model, a novel framework that combines semantic and geometric graph Transformer for 3D Gaussian Splating-based Occupancy Prediction. We propose the Dual Gaussians Graph Attenntion, which dynamically constructs dual graph structures: a geometric graph adaptively calculating KNN search radii based on Gaussian poses, enabling large-scale Gaussians to aggregate features from broader neighborhoods while compact Gaussians focus on local geometric consistency; a semantic graph retaining top-M highly correlated nodes via cosine similarity to explicitly encode semantic relationships within and across instances. Coupled with the Multi-scale Graph Attention framework, fine-grained attention at lower layers optimizes boundary details, while coarse-grained attention at higher layers models object-level topology. Experiments on the SurroundOcc dataset achieve an mIoU of 24.10%, reducing GPU memory to 6.1 GB, demonstrating a 1.97% mIoU improvement and 13.7% memory reduction compared to GaussianWorld.
@@ 1 Introduction
In recent years, there has been a notable transition in the field of autonomous driving technology, with a shift from LiDAR-based multi-modal fusion to image-based vision-centric perception. This transition has been driven primarily by the potential to reduce reliance on costly LiDAR sensors [1, 2]. The occupancy network has emerged as a cornerstone in vision-centric methodologies to capture the dense 3D structure of the real world [3, 4]. This developing field of perception technology aims to infer the occupied state of each voxel in a voxelized world instead of predicting 3D box [5, 6, 7] or vectorized mapping [8, 9, 10, 11, 12]. It demonstrates a robust capacity for generalization, enabling its application to open-set objects, irregularly shaped vehicles and specialized road structures [13, 14].
@@
3D semantic occupancy prediction, as a dense volumetric segmentation task [16, 17], fundamentally requires efficient yet expressive scene representations. While voxel-based approaches [18, 15] employ dense cubic partitions to capture geometric details with high fidelity, their computational demands scale cubically with spatial resolution due to inherent redundancy in empty space representation. To address this limitation, planar projection techniques such as bird’s-eye-view (BEV) [19, 20] and trilinear positional value (TPV) [21] have been proposed, which collapse volumetric grids along specific axes to form compact 2D feature maps. However, these dimensionality reduction strategies inevitably retain unnecessary spatial regions in their feature encoding, thereby constraining model efficiency and representational capacity. In contrast to these spatially continuous paradigms, recent object-centric frameworks based on 3D Gaussian Splating (3DGS) [22] introduce a paradigm shift through probabilistic spatial decomposition, where each Gaussian component parametrizes local geometry (via mean and covariance), visibility (opacity), and semantic attributes in a learnable, sparse formulation. This statistical representation effectively decouples scene complexity from computational cost while preserving semantic expressiveness. However, several limitations persist in the current 3DGS-based methods: (1) Unified feature aggregation neglects semantic correlations across distant regions, leading to contextual fragmentation. (2) Iterative refinement via MLPs lacks explicit geometric constraints, causing position drift and semantic ambiguities at object boundaries.
To address these challenges, we propose GraphGSOcc, a dynamic structure-aware model that injects geometric and semantic priors into Gaussian interactions. Specifically, we dynamically construct dual graph structures. We implement adaptive radius calculation for KNN search in the geometric graph, where the search range for each Gaussian is dynamically scaled based on its poses, enabling large-scale Gaussians (e.g., road surfaces) to aggregate features from broader neighborhoods while compact Gaussians (e.g., pedestrians) focus on local geometric consistency. For the semantic graph, we compute feature affinity matrices across all Gaussians and selectively retain top-M nodes with highest cosine similarity, explicitly encoding cross-instance and similar instance relationships. Further, a multi-scale aggregation framework hierarchically refines Gaussians—fine-grained attention at lower layers optimizes covariance matrices for boundary details (e.g., small objects prediction), while coarse-grained attention at higher layers models object-level topology (e.g., pedestrian-vehicle motion patterns).
@@
    Our method achieves a mIoU of 24.10% on Surroundocc [15], while also reducing the GPU memory usage to 6.1 GB, as shown in Fig. 1.
@@ 2 Related Work
2.2 Grid-based 3D Semantic Occupancy Prediction
Recent advances in scene perception for autonomous driving have seen plane representations emerge as a competitive approach, offering compelling alternatives to traditional LiDAR-based methods. Among these, BEVFormer [19] pioneered the use of purely camera inputs while achieving comparable performance in detection and segmentation tasks. This method transforms image features into bird’s-eye-view (BEV) [28, 29, 30, 31, 32] features as a unified scene representation, capitalizing on the viewpoint’s information diversity for downstream applications. However, the BEV paradigm faces limitations in 3D occupancy prediction due to its inherent loss of height information [15]. Addressing this, TPVFormer [21] introduced an enhanced tri-perspective view representation that explicitly incorporates height dimensions, better accommodating 3D scene understanding. Parallel developments have explored voxel-based representations [18, 15, 33, 24, 34] as a more precise alternative for 3D volumetric semantic prediction, offering finer granularity. Yet these grid-based approaches maintain uniform treatment of all voxels regardless of environmental sparsity, leading to fundamental inefficiencies in representation.
@@ 2.2 Grid-based 3D Semantic Occupancy Prediction
2.3 3D Gaussian Splatting for 3D Occupancy Prediction
Our proposed GraphGSOcc diverges from prior works in two works: (1) By dynamically constructing the graph structure between Gaussian distributions, the Dual Gaussian Graph Attention (DGGA) enhances local geometric and semantic correlations while preserving global modeling capabilities. (2) A Multi-scale Graph Attention (MGA) module is proposed to enhance the geometric detail prediction of small objects.
@@ 3 Proposed Method
In this section, we outline how the proposed DSOcc is utilized to enhance the generative power and efficiency of model facing feature loss and occlusion for 3D semantic occupancy prediction task. In the following sections, we first provide a brief introduction of the overall framework in Section 3.1. Thereafter, we provide a detailed explanation of Dual Gaussians Graph Attention (DGGA) which is the core component of GraphGSOcc in Section 3.2. Additionally, the Multi-scale Graph Attention (MGA) layers are demonstrated in details in Section 3.3. Finally, we introduce the other components of GraphGSOcc in Section 3.4.
@@ 3.1 Framework Overview
Fig. 2 depicts the pipeline of our GraphGSOcc. As denoted in Fig. 2, the input data for GraphGSOcc consists of surround-view images sequence, while the output is 3D semantic occupancy prediction results. At current frame $t$, surround-view images are first processed by the 2D Encoder. Then, we load the historical frame $t-1$ gaussians and initialize the random prior gaussians. We use Temporal Gaussians Align based on GaussianWorld to get the aligned gaussians and complete gaussians at current frame $t$. Additionally, we use Self-Encoding and Cross-Attention to optimize the gaussians based on the image features. Furthermore, the two key components: (i) Dual Gaussian Graph Attention (DGGA) and (ii) Multi-scale Graph Attention (MGA) are introduced to enhance and refine the geometric and semantic correlations between gaussians by the means and the feature of gaussians. Finally, the enhanced gaussians are used to generate the 3D voccupancy at current frame $t$ via GSHead.
@@ 3.2 Dual Gaussian Graph Attention
In this section, as shown in Figs. 3, we explain (i) how we build the geometric graph structure and semantic graph structure and (ii) how we aggregate the features of neighbor nodes to the center nodes and (iii) how we fuse the features of geometric graph and semantic graph.
@@ 3.2.3 Adaptive Fusion
3.3 Multi-scale Graph Attention
In this section, we elaborate on the design of a Multi-scale Graph Attention (MGA) mechanism, which is tailored to fuse 3D Gaussians with diverse Top-K and Top-M configurations for hierarchical feature aggregation. As illustrated in Fig. 5, the mechanism constructs multi-scale graph structures by defining dynamic neighbor node ranges, specifically setting Top-K as [100, 75, 50, 20] and Top-M as [100, 75, 50, 20]. This strategic parameterization enables the capture of contextual information at varying spatial scales—larger Top values (e.g., 100) facilitate global contextual reasoning, while smaller values (e.g., 20) emphasize local geometric details.Based on these graphs, we generate 3D Gaussians through the DGGA module under different scale graph structures. Finally, we fuse multiple 3D Gaussians using Adaptive Fusion to obtain the final output 3D Gaussian $G_{final}$.

### v2

# GraphGSOcc: Semantic-Geometric Graph Transformer with Dynamic-Static Decoupling for 3D Gaussian Splatting-based Occupancy Prediction

@@ Abstract
Addressing the task of 3D semantic occupancy prediction for autonomous driving, we tackle two key issues in existing 3D Gaussian Splatting (3DGS) methods: (1) unified feature aggregation neglecting semantic correlations among similar categories and across regions, (2) boundary ambiguities caused by the lack of geometric constraints in MLP iterative optimization and (3) biased issues in dynamic-static object coupling optimization. We propose the GraphGSOcc model, a novel framework that combines semantic and geometric graph Transformer and decouples dynamic-static objects optimization for 3D Gaussian Splatting-based Occupancy Prediction. We propose the Dual Gaussians Graph Attenntion, which dynamically constructs dual graph structures: a geometric graph adaptively calculating KNN search radii based on Gaussian poses, enabling large-scale Gaussians to aggregate features from broader neighborhoods while compact Gaussians focus on local geometric consistency; a semantic graph retaining top-M highly correlated nodes via cosine similarity to explicitly encode semantic relationships within and across instances. Coupled with the Multi-scale Graph Attention framework, fine-grained attention at lower layers optimizes boundary details, while coarse-grained attention at higher layers models object-level topology. On the other hand, we decouple dynamic and static objects by leveraging semantic probability distributions and design a Dynamic-Static Decoupled Gaussian Attention mechanism to optimize the prediction performance for both dynamic objects and static scenes. GraphGSOcc achieves state-ofthe-art performance on the SurroundOcc-nuScenes, Occ3D-nuScenes, OpenOcc and KITTI occupancy benchmarks. Experiments on the SurroundOcc dataset achieve an mIoU of 25.20%, reducing GPU memory to 6.8 GB, demonstrating a 1.97% mIoU improvement and 13.7% memory reduction compared to GaussianWorld.
@@ 1 Introduction
Recent years have witnessed a significant paradigm shift in autonomous driving perception, transitioning from multi-modal fusion reliant on LiDAR toward vision-centric approaches based predominantly on image data. This evolution is primarily motivated by the potential to reduce dependence on costly LiDAR sensors [1]. Within this vision-centric framework, occupancy networks have become foundational, capturing dense 3D structures of the environment [2]. This burgeoning perception paradigm focuses on inferring the occupancy state of individual voxels within a discretized 3D space, representing a departure from methods predicting 3D bounding boxes [3, 4, 5, 6] or vectorized map elements [7, 8, 9, 10, 11]. Occupancy networks demonstrate strong generalization capabilities, effectively handling open-set objects, irregularly shaped vehicles, and specialized road infrastructure [12, 13, 14].
@@
3D semantic occupancy prediction, as a dense volumetric segmentation task [16, 17], fundamentally requires efficient yet expressive scene representations. While voxel-based approaches [18, 15] employ dense cubic partitions to capture geometric details with high fidelity, their computational demands scale cubically with spatial resolution due to inherent redundancy in empty space representation. To address this limitation, planar projection techniques such as bird’s-eye-view (BEV) [19, 20] and trilinear positional value (TPV) [21] have been proposed, which collapse volumetric grids along specific axes to form compact 2D feature maps. However, these dimensionality reduction strategies inevitably retain unnecessary spatial regions in their feature encoding, thereby constraining model efficiency and representational capacity. In contrast to these spatially continuous paradigms, recent object-centric frameworks based on 3D Gaussian Splatting (3DGS) [22] introduce a paradigm shift through probabilistic spatial decomposition, where each Gaussian component parametrizes local geometry (via mean and covariance), visibility (opacity), and semantic attributes in a learnable, sparse formulation. This statistical representation effectively decouples scene complexity from computational cost while preserving semantic expressiveness. However, several limitations persist in the current 3DGS-based methods: (1) Unified feature aggregation neglecting semantic correlations among similar categories and across regions, leading to contextual fragmentation. (2) Iterative refinement via MLPs lacks explicit geometric constraints, causing position drift and semantic ambiguities at object boundaries. (3) Biased issues in dynamic-static object coupling optimization.
To address these challenges, we propose GraphGSOcc, a dynamic structure-aware model that injects geometric and semantic priors into Gaussian interactions. Specifically, we dynamically construct dual graph structures. We implement adaptive radius calculation for KNN search in the geometric graph, where the search range for each Gaussian is dynamically scaled based on its poses, enabling large-scale Gaussians (e.g., road surfaces) to aggregate features from broader neighborhoods while compact Gaussians (e.g., pedestrians) focus on local geometric consistency. For the semantic graph, we compute feature affinity matrices across all Gaussians and selectively retain top-M nodes with highest cosine similarity, explicitly encoding cross-instance and similar instance relationships. Further, a multi-scale aggregation framework hierarchically refines Gaussians—fine-grained attention at lower layers optimizes covariance matrices for boundary details (e.g., small objects prediction), while coarse-grained attention at higher layers models object-level topology (e.g., pedestrian-vehicle motion patterns). Finally, we decouple dynamic and static objects by leveraging semantic probability distributions and design a Dynamic-Static Decoupled Gaussian Attention mechanism to optimize the prediction performance for both dynamic objects and static scenes.
@@
    We propose a Dynamic-Static Decoupled Gaussian Attention (DSDGA) mechanism to optimize the prediction for dynamic and static objects.
  • $\bullet$
    Our method achieves a mIoU of 25.20% on Surroundocc [15], while also reducing the GPU memory usage to 6.8 GB, as shown in Fig. 1. Our GraphGSOcc also achieves state-of-the-art performance on OpenOcc [23], Occ3D [17] and SSCBench-KITTI-360 [24].
@@ 2 Related Work
2.2 Scene-based 3D Semantic Occupancy Prediction
In recent advancements of scene perception for autonomous driving, plane representations have emerged as a competitive approach, presenting compelling alternatives to conventional LiDAR-based methods. Among these, BEVFormer [19] was the first to utilize purely camera inputs while achieving comparable performance in detection and segmentation tasks. This method transforms image features into bird’s-eye-view (BEV) [30, 31, 32, 33, 34] features as a unified scene representation, leveraging the viewpoint’s information diversity for downstream applications. Nevertheless, the BEV paradigm encounters constraints in 3D occupancy prediction because of its inherent loss of height information [15]. To address this, TPVFormer [21] introduced an enhanced tri-perspective view representation that explicitly integrates height dimensions, better suiting 3D scene understanding. Concurrently, parallel developments have explored voxel-based representations [18, 15, 35, 26, 36] as a more precise option for 3D volumetric semantic prediction, providing finer granularity. However, these grid-based approaches uniformly handle all voxels without considering environmental sparsity, leading to inherent inefficiencies in representation.
2.3 3D Gaussian Splatting for 3D Occupancy Prediction
Our proposed GraphGSOcc diverges from prior works in three works: (1) By dynamically constructing the graph structure between Gaussian distributions, the Dual Gaussian Graph Attention (DGGA) enhances local geometric and semantic correlations while preserving global modeling capabilities. (2) A Multi-scale Graph Attention (MGA) module is proposed to enhance the geometric detail prediction of small objects. (3) We decouple dynamic and static objects by leveraging semantic probability distributions and design a Dynamic-Static Decoupled Gaussian Attention mechanism to optimize the prediction performance for both dynamic objects and static scenes.
@@ 2.2 Scene-based 3D Semantic Occupancy Prediction
Figure 6: An overview of Dynamic-Static Decoupled Gaussian Attention (DSDGA) framework,
@@ 3 Proposed Method
In this section, we outline how the proposed GraphGSOcc is utilized to enhance the performance for 3D semantic occupancy prediction task. In the following sections, we first provide a brief introduction of the overall framework in Section 3.1. Thereafter, we provide a detailed explanation of Dual Gaussians Graph Attention (DGGA) which is the core component of GraphGSOcc in Section 3.2. Additionally, the Dynamic-Static Decoupled Gaussian Attention (DSDGA) layer, which is the other core component of GraphGSOcc, is demonstrated in details in Section 3.3. Finally, we introduce the other components of GraphGSOcc in Section 3.4.
@@ 3.1 Framework Overview
Fig. 2 depicts the pipeline of our GraphGSOcc. As denoted in Fig. 2, the input data for GraphGSOcc consists of surround-view images sequence, while the output is 3D semantic occupancy prediction results. At current frame $t$, surround-view images are first processed by the 2D Encoder. Then, we load the historical frame $t-1$ gaussians and initialize the random prior gaussians. We use Temporal Gaussians Align based on GaussianWorld to get the aligned gaussians and complete gaussians at current frame $t$. Additionally, we use Self-Encoding and Cross-Attention to optimize the gaussians based on the image features. Furthermore, the two key components: (i) Dual Gaussian Graph Attention (DGGA) is introduced to enhance and refine the geometric and semantic correlations between gaussians by the means and the feature of gaussians, and (ii) Dynamic-Static Decoupled Gaussian Attention (DSDGA) is introduced to decouple the dynamic and static gaussians. Finally, the enhanced gaussians are used to generate the 3D occupancy at current frame $t$ via GSHead.
@@ 3.2 Dual Gaussian Graph Attention
In this section, as shown in Figs. 3, we explain (i) how we build the geometric graph structure and semantic graph structure and (ii) how we aggregate the features of neighbor nodes to the center nodes and (iii) how we fuse the features of geometric graph and semantic graph and (iv) how we adaptively fuse the multi-scale gaussian features.
@@ 3.2.2 Geometric and Semantic Graph Attention
Figure 8: Visual comparison of our GraphGSOcc with state-of-the-art methods on Occ3D. Compared to GaussianFormer, our GraphGSOcc generates more precise prediction results of semantic classes and complex scenarios (labeled in red circles).
Table 1: Performance comparison of 3D Occupancy Prediction Methods on the NuScenes SurroundOcc Dataset. The Latency and Memory Consumption for All Methods are Tested on One NVIDIA 4090 GPU with Batchsize One. The Bold and Underline Indicate the Best and Second-best Tesults, respectively. (TPVFormer* is Supervised by Dense Occupancy Annotations. GraphGSOcc** Denotes Ours Method Based on GaussianFormer. GraphGSOcc* Denotes Ours Method Based on GaussianFormer2. GraphGSOcc Denotes Our Method Based on GaussianWorld)
MonoScene CVPR’22 ResNet-101 1600 $\times$ 928 7.31 23.96 - -
BEVFormer ECCV’22 ResNet-101 1600 $\times$ 900 16.75 30.50 212 6651
TPVFormer CVPR’23 ResNet-101 1600 $\times$ 928 11.66 11.51 341 6926
TPVFormer* CVPR’23 ResNet-101 1600 $\times$ 928 17.10 30.86 341 6926
OccFormer ICCV’23 ResNet-101 1600 $\times$ 928 19.03 31.39 - -
SurroundOcc ICCV’23 ResNet-101 1600 $\times$ 900 20.30 31.49 312 5491
GraphGSOcc** - ResNet-101 1600 $\times$ 864 22.41 34.21 280 4120
GraphGSOcc* - ResNet-101 1600 $\times$ 864 23.75 35.58 270 2812
GraphGSOcc - ResNet-101 1600 $\times$ 864 25.20 36.11 368 6063
Table 2: 3D Occupancy Prediction Performence on OpenOcc and Occ3D Dataset. The RayIoU Results are Obtained Using the Annotations from OpenOcc [23], While the MIoU Results are Based on the Occ3D Annotations [17]. Bold and Underline Denote the Best Performence and the Second-Best Performence.
Methods Backbone Input Size RayIoU mIoU
RenderOcc R101 1600$\times$900 19.5 24.6
OccNet R101 1600$\times$900 32.7 36.8
BEVFormer R101 1600$\times$900 28.1 39.1
FB-Occ R101 1600$\times$900 33.5 39.4
SparseOcc R101 1600$\times$900 35.1 30.6
GraphGSOcc R101 1600$\times$900 36.7 42.6
Table 3: Results on the SSCBench-KITTI-360 Test Set [24] with a Monocular Camera. GraphGSOcc Achieves New State-of-the-art, Achieving Strong Performance in MIoU and IoU
Methods Backbone Input IoU mIoU
MonoScene R101 C 37.87 12.31
Voxformer R101 C 18.76 11.91
TPVFormer R101 C 40.22 13.64
OccFormer R101 C 40.27 13.87
GaussianFormer R101 C 35.38 12.92
GaussianFormer-2 R101 C 38.37 13.90
GraphGSOcc R101 C 40.92 15.58
Table 4: Ablation on the Proposed DGGA Layer without DSDGA. Dyna. Object Denotes Dynamic Object. Stat. Object Denotes Static Object.
Methods mIoU IoU
All Dyna. object Stat. object All
baseline 18.41 16.73 20.09 29.23
w/ MLP 19.10 16.99 21.21 29.83
w/ GGA 19.83 17.21 22.45 31.98
w/ SGA 20.52 18.33 22.71 30.21
w/ DGGA 21.01 18.54 23.48 32.32
Table 5: Ablation on the MGA module without DSDGA.
Methods mIoU IoU Memory (MB)
Add 19.52 31.25 3080
Concat 19.48 31.10 3288
Conv Fusion 20.21 31.58 3250
Adaptive Fusion 21.31 33.41 3320
Figure 12: Demonstrating the effect of the DSDGA in the "pedestrian" dynamic objects. The first column is the model with DSDGA module. The second column is the model without DSDGA module. The first row is the GT. The second row is the prediction occupancy. The third row is the heatmap.
Table 6: Ablation on the parameters of Top-K and Top-M module without DSDGA.
Methods Top-K Top-M Memory (MB) mIoU IoU
A [100, 100, 100, 100] [100, 100, 100, 100] 3500 21. 22 33.22
B [200, 150, 100, 50] [200, 150, 100, 50] 3865 21.12 32.98
C [50, 25, 10, 5] [50, 25, 10, 5] 2010 19.51 31.21
D [100, 75, 50, 20] [100, 75, 50, 20] 3320 21.31 33.41
Table 7: Ablation on the DSDGA Module without DGGA and MGA.
Methods mIoU IoU
All Dyna. object Stat. object All
baseline 18.41 16.73 20.09 29.23
w/ DCA 19.85 18.99 20.71 30.93
w/ SCA 20.06 16.99 23.13 31.88
w/ DSDGA 22.38 19.89 24.88 32.95
@@ 3.2.3 Adaptive Fusion
3.2.4 Multi-scale Graph Attention
In this section, we elaborate on the design of a Multi-scale Graph Attention (MGA) mechanism, which is tailored to fuse 3D Gaussians with diverse Top-K and Top-M configurations for hierarchical feature aggregation. As illustrated in Fig. 5, the mechanism constructs multi-scale graph structures by defining dynamic neighbor node ranges, specifically setting Top-K as [100, 75, 50, 20] and Top-M as [100, 75, 50, 20]. This strategic parameterization enables the capture of contextual information at varying spatial scales—larger Top values (e.g., 100) facilitate global contextual reasoning, while smaller values (e.g., 20) emphasize local geometric details. Based on these graphs, we generate 3D Gaussians through the DGGA module under different scale graph structures. Finally, we fuse multiple 3D Gaussians using Adaptive Fusion to obtain the multi-scale output 3D Gaussian $G_{multiscale}$.
3.3 Dynamic-Static Decoupled Gaussian Attention
In this section, we primarily introduce the Dynamic-Static Decoupled Gaussian Attention (DSDGA) module to further optimize the prediction for dynamic and static objects. This module achieves fine-grained feature separation by leveraging semantic information to distinguish Gaussian anchors into dynamic and static groups, and employs decouple cross attention mechanisms to enable bidirectional feature interaction. Firstly, we decouple dynamic and static Gaussians based on the semantic masks:
$S=\text{softplus}(G_{:,s:s+d}),$ (7)
$c_{i}=\arg\max_{j=1}^{d}S_{ij},\quad i=1,2,\ldots,N,$
$m_{\text{static}}[i]=\begin{cases}1,&\text{if}\text{ }c_{i}\geq t,\\ 0,&\text{otherwise},\end{cases}$
$m_{\text{dynamic}}=\neg m_{\text{static}},$
where $G$ is the 3D Gaussian, $S$ is the semantic score matrix, $c_{i}$ is the index of the maximum semantic score for the $i$-th Gaussian, $m_{\text{static}}[i]$ is the static mask for the $i$-th Gaussian, $m_{\text{dynamic}}[i]$ is the dynamic mask for the $i$-th Gaussian.
$G_{\text{dynamic}}=G\odot m_{\text{dynamic}},$ (8)
$G_{\text{static}}=G\odot m_{\text{static}},$
where $G_{\text{dynamic}}$ is the dynamic 3D Gaussian, and $G_{\text{static}}$ is the static 3D Gaussian. Subsequently, after obtaining the decoupled dynamic object Gaussians and static scene Gaussians, we employ the proposed Dynamic-Static Decoupled Gaussian Attention (DSDGA) module to interactively optimize the representations of both dynamic objects and static scenes. This module facilitates bidirectional feature interaction through cross-scale attention mechanisms, enabling the exchange of complementary information between the dynamic and static components. Specifically, the DSDGA module consists of two key sub-components:
3.3.1 Dynamic Cross Attention (DCA)
This sub-module refines the dynamic object representations by leveraging structural and contextual information from the static scene. For instance, it uses prior knowledge of static scene layouts to improve motion prediction of dynamic objects:
$G^{{}^{\prime}}_{dynamic}=DCA(Q_{G_{dynamic}},K_{G_{static}},V_{G_{static}})$ (9)
3.3.2 Static Cross Attention (SCA)
This sub-module enhances the static scene representation by selectively attending to dynamic object features that are relevant to scene understanding. For example, it captures how moving agents interact with static environments (e.g., pedestrians on sidewalks, vehicles on roads):
$G^{{}^{\prime}}_{static}=SCA(Q_{G_{static}},K_{G^{{}^{\prime}}_{dynamic}},V_{G^{{}^{\prime}}_{dynamic}})$ (10)
Finally, we fuse the refined dynamic and static Gaussian representations $G^{{}^{\prime}}_{\text{dynamic}}$ and $G^{{}^{\prime}}_{\text{static}}$ through concatenation to generate the final output $G_{\text{final}}$.